A projectile is fired in such a way that its horizontal range is equal to three times its maximum height. What is the angle of projection?


1
Expert's answer
2021-08-16T08:35:37-0400

The range:


"R=\frac{v^2\sin(2\theta)}{g}."

Height:


"H=\frac{v^2\sin^2(\theta)}{2g}."

We know that


"R=3H,\\\space\\
\frac{v^2\sin(2\theta)}{g}=3\frac{v^2\sin^2(\theta)}{2g},\\\space\\
2\sin\theta\cos\theta=\frac32\sin^2\theta,\\\space\\
\tan\theta=\frac43,\\\space\\
\theta=53°."